# Onkar Susladkar1, Jishu Sen Gupta2, Chirag Sehgal3, Sparsh Mittal4, Rekha Singhal5

### v1

@@ Abstract
The spatio-temporal complexity of video data presents significant challenges in tasks such as compression, generation, and inpainting. We present four key contributions to address the challenges of spatiotemporal video processing. First, we introduce the 3D Mobile Inverted Vector-Quantization Variational Autoencoder (3D-MBQ-VAE), which combines Variational Autoencoders (VAEs) with masked token modeling to enhance spatiotemporal video compression. The model achieves superior temporal consistency and state-of-the-art (SOTA) reconstruction quality by employing a novel training strategy with full frame masking. Second, we present MotionAura, a text-to-video generation framework that utilizes vector-quantized diffusion models to discretize the latent space and capture complex motion dynamics, producing temporally coherent videos aligned with text prompts. Third, we propose a spectral transformer-based denoising network that processes video data in the frequency domain using the Fourier Transform. This method effectively captures global context and long-range dependencies for high-quality video generation and denoising. Lastly, we introduce a downstream task of Sketch Guided Video Inpainting. This task leverages Low-Rank Adaptation (LoRA) for parameter-efficient fine-tuning. Our models achieve SOTA performance on a range of benchmarks. Our work offers robust frameworks for spatiotemporal modeling and user-driven video content manipulation. We will release the code, datasets, and models in open-source.
@@ 1 Introduction
Video generation refers to generating coherent and realistic sequences of frames over time, matching the consistency and appearance of real-world videos. Recent years have witnessed significant advances in video generation based on learning frameworks ranging from generative adversarial networks (GANs) (Aldausari et al., 2022), diffusion models (Ho et al., 2022a), to causal autoregressive models (Gao et al., 2024; Tudosiu et al., 2020; Yu et al., 2023b). Video generative models facilitate a range of downstream tasks such as class conditional generation, frame prediction or interpolation (Ming et al., 2024), conditional video inpainting (Zhang et al., 2023; Wu et al., 2024) and video outpainting (Dehan et al., 2022). Video generation has several business use cases, such as marketing, advertising, entertainment, and personalized training.
@@
To address the challenges of visual content generation, we propose MotionAura. At the core of MotionAura is our novel VAE, named 3D-MBQ-VAE (3D MoBile Inverted VQ-VAE), which achieves good temporal comprehension for spatiotemporal compression of videos. For the video generation phase, we use a denoiser network comprising a series of proposed Spectral Transformer blocks trained using the masked token modeling approach. Our novelty lies both in the architectural changes in our transformer blocks and the training pipeline. These include using a 2D Fast Fourier Transform (FFT) layer and Rotary Positional Embeddings (Su et al., 2021). Following an efficient and robust, non-autoregressive approach, the transformer block predicts all the masked tokens at once. To ensure temporal consistency across the video latents, during tokenization, we randomly mask one of the frames completely and predict the index of the masked frame. This helps the model learn the frame sequence and improve the temporal consistency of frames at inference time. Finally, we address a downstream task of sketch-guided video inpainting. To our knowledge, ours is the first work to address the sketch-guided video inpainting task. Using our pre-trained noise predictor as a base and the LORA adaptors (Hu et al., 2021), our model can be finetuned for various downstream tasks such as class conditional video inpainting and video outpainting. Figure [Uncaptioned image]  MotionAura: Generating High-Quality and Motion Consistent Videos using Discrete Diffusion shows examples of text-conditioned video generation using MotionAura. Our contributions are:
@@ 2 Related Works on visual content generation
Image generation: Text-guided image synthesis has resulted in high quality and spatially coherent data (Esser et al., 2024; Sun et al., 2024). Generative Adversarial Networks (Sauer et al., 2022; Karras et al., 2020; 2019; 2021; Goodfellow et al., 2014) model visual distributions to produce perceptually rich images but fail to capture the complete spectrum of data distribution. Moreover, these networks are difficult to optimize. Kingma & Welling (2019) introduce another approach to generating visual data wherein the decoders of the autoencoder-like architectures are trained to reconstruct images by sampling over the distributions provided by encoders. This regularization strategy helps to learn continuous and meaningful latent space. Quantization of the produced latents (van den Oord et al., 2018) further improves the quality of generated data using various sampling methods over the learned latents. While existing sampling methods can capture dense representations in their latent space, they fail to produce sharp, high-quality images. Diffusion models train networks to learn a denoising process, appropriately fitting the inductive bias in image data. Recent developments such as DDPMs (Ho et al., 2020), latent diffusion methods (Rombach et al., 2022a) and diffusion transformers (Esser et al., 2024; Peebles & Xie, 2023; Chen et al., 2024; 2023) have revolutionized the field of image generation.
Video generation: Video generation is even more challenging than image generation since generating high-quality videos demands both spatial and temporal consistencies. For video generation, researchers have used both transformer models and diffusion models. VideoGPT (Yan et al., 2021) generates videos autoregressively using the latents obtained from the VQ-VAE tokenizer. Processing videos in the latent space rather than the pixel space reduces training and inference latency. The early diffusion-based video generation models transferred weights from the TextToImage (T2I) approaches and adapted them by inflating the convolutional backbone to incorporate the temporal dynamics (Ho et al., 2022b). Latent diffusion models (LDM) based Text2Video (T2V) approaches generate higher-quality videos of longer duration (Blattmann et al., 2023). However, diffusion-based approaches using convolution backbones fail to produce high frame-rate videos and generally rely on other architectures to interpolate intermediate frames. Recent works have combined transformer-based predictions with diffusion-based pretraining (Yu et al., 2023a; 2024; Kondratyuk et al., 2024).
@@ 3.1 Pre-Training of 3D-MBQ-VAE
We employed a 3D-VQ-VAE (Vector Quantized Variational Autoencoder) to tokenize videos in 3D space; refer Appendix B for the architecture of 3D-MBQ-VAE. We pretrained our network on the YouTube-100M dataset to ensure that it produces video-appropriate tokens. The extensive content variety in the YouTube-100M dataset enhances the model’s generalization and representational capabilities. Figure 2 shows our pretraining approach for learning efficient latents with temporal and spatial consistencies. The pretraining leverages two approaches for optimal video compression and discretization: (i) Random masking of frames (He et al., 2021) using 16$\times$16 and 32$\times$32 patches. This training method ensures self-supervised training of the autoencoder to efficiently learn spatial information. (ii) Complete masking of a single frame in the sequence to enforce learning of temporal information. Let $B$ be the batch size, $C$ be number of channels, $N$ the number of frames per video, $H$, and $W$ be the height and width of the video frames, respectively. For a given video $V\in\mathbb{R}^{(B,N,3,H,W)}$, a randomly selected frame $f_{i}\in\mathbb{R}^{(3,H,W)}$ is masked completely. The encoder $\mathcal{E}(V)$ returns $z_{c}\in\mathbb{R}^{(B,C,N,H/16,W/16)}$. This compressed latent in the continuous domain is fed into the MLP layer to return $P(f_{i})$, which represents the probability of $i^{th}$ frame being completely masked.
@@ 3.3 Proposed Spectral Transformer
As shown in Figure 4, before computing the self-attention, the latent representation ${B}_{t}$ is transformed into the frequency domain using a 2D Fourier Transform Layer (fft2D) for each of $\mathbf{K}$, $\mathbf{V}$, and $\mathbf{Q}$. This layer applies two 1D Fourier transforms - along the embedding dimension and the sequence dimension thereby converting spatial domain data into the frequency domain. This enables the segregation of high-frequency features from low-frequency ones, enabling more efficient manipulation and learning of spatial features. We take only the real part of fft2D layer outputs to make all quantities differentiable (Eq. 2). Here, MSA stands for multi-headed self-attention.
@@
Finally, an MLP followed by a softmax layer gives us the predicted denoised latent given $z_{t}$. This process is repeated until we get the completely denoised latent $P(z_{0}\mid z_{t},t,T(X_{t}))$ Finally, these output latents are mapped back into the video domain using our pre-trained 3D-MBQ-VAE decoder $\mathcal{D}$.
@@ 5 Experimental results
We now present experimental results of i) 3D-MBQ-VAE ii) MotionAura for text-guided video generation and iii) MotionAura for sketch-guided video inpainting. The details of experimental setup are provided in Appendices C and D. Additional qualitative results are provided in Appendices A, E and F.
@@ 5.1 Latent Reconstruction Results of 3D-MBQ-VAE
For evaluating our 3D-MBQ-VAE, we selected the COCO-2017 and the WebVID validation datasets. Following Zhao et al. (2024), we crop each frame to 256$\times$256 resolution and sample 48 frames per video sequentially. As shown in Table 1, our proposed 3D-MBQ-VAE consistently outperforms SOTA 3D VAEs across all metrics. This can be attributed to a novel combination of training strategies and the enhanced loss function. Randomly masking parts of various frames He et al. (2021) enables our model to efficiently learn spatial information. Further, inclusion of $\mathcal{L}_{\text{MFI}}$ in the loss function emphasizes learning efficient temporal features.
Table 1: Quantitative comparison of VAEs for video reconstruction task on COCO-Val and WebVid-Val datasets. Frame compression rate (FCR) is the ratio between the size of video frame before and after compression. Compatibility (comp.) represents if the model can be used as a VAE for existing generative models. Our method demonstrates superior performance on all metrics.
@@
3D-MBQ-VAE (Ours) 140M + 177M 4$\times$ 31.2 / 0.848 / 0.092 32.1 / 0.858 / 0.108
@@
Figure 8: The videos corresponding to these frames show the improvement in motion quality brought by FFT layers. Click on the link to see the videos.
@@ 5.2 Results of Text Conditioned Video Generation
Table 3: Results of the text-conditioned video generation (Text2Video) models. Recaptioning the dataset improves all the metrics. Inference Time was calculated on a single A100.
Method Params WebVID 10M WebVid 10M - recaptioned Inf Tims(s)
@@ 6 Conclusion
B.1 Encoder
The encoder takes an input video with dimensions $B\times C\times N\times H\times W$, where $B$ is the batch size, $C$ is the number of channels, and $N\times H\times W$ represents the spatiotemporal dimensions (i.e., frames, height, and width). The encoding begins with a Conv3D block that captures spatiotemporal features using a 3D convolution with a kernel size of $1\times 3\times 3$, followed by a stride of $1\times 2\times 2$ to reduce spatial resolution while expanding the channel depth (e.g., from 3 to 64 channels). The encoder then applies multiple 3D-MB Conv blocks (Mobile Inverted Conv Blocks) to capture complex features using depthwise separable convolutions, maintaining computational efficiency. These blocks are interspersed with Inception-Fused-MB Conv blocks, which improve feature extraction by combining multiple filter sizes and dimensions, enabling multi-scale processing.
@@ B.3 Key Features
B.4 Mobile-Inverted 3D Conv Block
@@
B.5 Inception-Fused MB Conv Block
@@
The 3D Inception-Fused-MB Conv Block is a sophisticated convolutional module designed to efficiently capture both spatial and temporal features in video data, commonly used in tasks like Video VQ-VAE (Vector Quantized Variational Autoencoder) shown in Figure S.8. This block employs an Inception-style structure with three parallel 3D convolution paths, each with different kernel configurations (e.g., 3 x 3 x 1, 3 x 1 x 3, and 1 x 3 x 3) to capture multi-scale spatiotemporal patterns in various dimensions. Each branch is followed by a depthwise separable 3D-MobileBlock (MB-Conv) that reduces computational complexity while maintaining efficiency. The outputs of these paths are concatenated, allowing the network to fuse multi-scale features, which are then processed by a final 3D convolutional block. This design enables efficient handling of complex video representations in VQ-VAEs, optimizing the model’s ability to compress and reconstruct video sequences while preserving key spatial and temporal information.

### v2

# Onkar Kishor Susladkar1, Jishu Sen Gupta2, Chirag Sehgal3, Sparsh Mittal4, Rekha Singhal5

@@ Abstract
The spatio-temporal complexity of video data presents significant challenges in tasks such as compression, generation, and inpainting. We present four key contributions to address the challenges of spatiotemporal video processing. First, we introduce the 3D Mobile Inverted Vector-Quantization Variational Autoencoder (3D-MBQ-VAE), which combines Variational Autoencoders (VAEs) with masked token modeling to enhance spatiotemporal video compression. The model achieves superior temporal consistency and state-of-the-art (SOTA) reconstruction quality by employing a novel training strategy with full frame masking. Second, we present MotionAura, a text-to-video generation framework that utilizes vector-quantized diffusion models to discretize the latent space and capture complex motion dynamics, producing temporally coherent videos aligned with text prompts. Third, we propose a spectral transformer-based denoising network that processes video data in the frequency domain using the Fourier Transform. This method effectively captures global context and long-range dependencies for high-quality video generation and denoising. Lastly, we introduce a downstream task of Sketch Guided Video Inpainting. This task leverages Low-Rank Adaptation (LoRA) for parameter-efficient fine-tuning. Our models achieve SOTA performance on a range of benchmarks. Our work offers robust frameworks for spatiotemporal modeling and user-driven video content manipulation. We release the code, datasets, and models in open-source link .
@@ 1 Introduction
Video generation refers to generating coherent and realistic sequences of frames over time, matching the consistency and appearance of real-world videos. Recent years have witnessed significant advances in video generation based on learning frameworks ranging from generative adversarial networks (GANs) (Aldausari et al., 2022), diffusion models (Ho et al., 2022a), to causal autoregressive models (Gao et al., 2024; Tudosiu et al., 2020; Yu et al., 2023c). Video generative models facilitate a range of downstream tasks such as class conditional generation, frame prediction or interpolation (Ming et al., 2024), conditional video inpainting (Zhang et al., 2023; Wu et al., 2024) and video outpainting (Dehan et al., 2022). Video generation has several business use cases, such as marketing, advertising, entertainment, and personalized training.
@@
To address the challenges of visual content generation, we propose MotionAura. At the core of MotionAura is our novel VAE, named 3D-MBQ-VAE (3D MoBile Inverted VQ-VAE), which achieves good temporal comprehension for spatiotemporal compression of videos. For the video generation phase, we use a denoiser network comprising a series of proposed Spectral Transformer blocks trained using the masked token modeling approach. Our novelty lies both in the architectural changes in our transformer blocks and the training pipeline. These include using a learnable 2D Fast Fourier Transform (FFT) layer and Rotary Positional Embeddings (Su et al., 2021). Following an efficient and robust, non-autoregressive approach, the transformer block predicts all the masked tokens at once. To ensure temporal consistency across the video latents, during tokenization, we randomly mask one of the frames completely and predict the index of the masked frame. This helps the model learn the frame sequence and improve the temporal consistency of frames at inference time. Finally, we address a downstream task of sketch-guided video inpainting. To our knowledge, ours is the first work to address the sketch-guided video inpainting task. Using our pre-trained noise predictor as a base and the LoRA adaptors (Hu et al., 2021), our model can be finetuned for various downstream tasks such as class conditional video inpainting and video outpainting. Figure [Uncaptioned image]  MotionAura: Generating High-Quality and Motion Consistent Videos using Discrete Diffusion shows examples of text-conditioned video generation using MotionAura. Our contributions are:
@@ 2 Related Works on visual content generation
Image generation: Text-guided image synthesis has resulted in high quality and spatially coherent data (Esser et al., 2024; Sun et al., 2024b). Generative Adversarial Networks (Sauer et al., 2022; Karras et al., 2021; Goodfellow et al., 2014) model visual distributions to produce perceptually rich images but fail to capture the complete spectrum of data distribution. Moreover, these networks are difficult to optimize. Kingma & Welling (2019) introduce another approach to generating visual data wherein the decoders of the autoencoder-like architectures are trained to reconstruct images by sampling over the distributions provided by encoders. This regularization strategy helps to learn continuous and meaningful latent space. Quantization of the produced latents (van den Oord et al., 2018; Susladkar et al., 2025) further improves the quality of generated data using various sampling methods over the learned latent. While existing sampling methods can capture dense representations in their latent space, they fail to produce sharp, high-quality images. Diffusion models train networks to learn a denoising process, appropriately fitting the inductive bias in image data. Recent developments such as DDPMs (Ho et al., 2020), latent diffusion methods (Rombach et al., 2022a) and diffusion transformers (Esser et al., 2024; Peebles & Xie, 2023; Chen et al., 2024; 2023) have revolutionized the field of image generation.
Video generation: Video generation is even more challenging than image generation since generating high-quality videos demands both spatial and temporal consistencies. For video generation, researchers have used both transformer models and diffusion models. VideoGPT (Yan et al., 2021) generates videos autoregressively using the latents obtained from the VQ-VAE tokenizer. Processing videos in the latent space rather than the pixel space reduces training and inference latency. The early diffusion-based video generation models transferred weights from the TextToImage (T2I) approaches and adapted them by inflating the convolutional backbone to incorporate the temporal dynamics (Ho et al., 2022b). Latent diffusion models (LDM) based Text2Video (T2V) approaches generate higher-quality videos of longer duration (Blattmann et al., 2023). However, diffusion-based approaches using convolution backbones fail to produce high frame-rate videos and generally rely on other architectures to interpolate intermediate frames. Recent works have combined transformer-based predictions with diffusion-based pretraining (Yu et al., 2023b; 2024; Kondratyuk et al., 2024).
@@ 3.1 Pre-Training of 3D-MBQ-VAE
We employed a 3D-VQ-VAE (Vector Quantized Variational Autoencoder) to tokenize videos in 3D space; refer Appendix B.1 for the architecture of 3D-MBQ-VAE. We pretrained our network on the YouTube-100M dataset to ensure that it produces video-appropriate tokens. The extensive content variety in the YouTube-100M dataset enhances the model’s generalization and representational capabilities. Figure 2 shows our pretraining approach for learning efficient latents with temporal and spatial consistencies. The pretraining leverages two approaches for optimal video compression and discretization: (i) Random masking of frames (He et al., 2021) using 16$\times$16 and 32$\times$32 patches. This training ensures self-supervised training of the autoencoder to efficiently learn spatial information. (ii) Complete masking of a single frame in the sequence to enforce learning of temporal information. Let $B$ be the batch size, $C$ be number of channels, $N$ the number of frames per video, $H$, and $W$ be the height and width of the video frames, respectively. For a given video $V\in\mathbb{R}^{(B,N,3,H,W)}$, a randomly selected frame $f_{i}\in\mathbb{R}^{(3,H,W)}$ is masked completely. The encoder $\mathcal{E}(V)$ returns $z_{c}\in\mathbb{R}^{(B,C,N,H/16,W/16)}$. This compressed latent in the continuous domain is fed into the MLP layer to return $P(f_{i})$, which shows the probability of $i^{th}$ frame being completely masked.
@@ 3.3 Proposed Spectral Transformer
As shown in Figure 4, before computing the self-attention, the latent representation ${B}_{t}$ is transformed into the frequency domain using a 2D Fourier Transform Layer (fft2D) for each of $\mathbf{K}$, $\mathbf{V}$, and $\mathbf{Q}$. This layer applies two 1D Fourier transforms (Susladkar et al., 2024; 2023) - along the embedding dimension and the sequence dimension thereby converting spatial domain data into the frequency domain. This enables the segregation of high-frequency features from low-frequency ones, enabling more efficient manipulation and learning of spatial features. We take only the real part of fft2D layer outputs to make all quantities differentiable (Eq. 2). Here, MSA stands for multi-headed self-attention.
@@
Finally, an MLP followed by a softmax layer gives the predicted denoised latent given $z_{t}$. This process is repeated until we get the completely denoised latent $P(z_{0}\mid z_{t},t,T(X_{t}))$ Finally, these output latents are mapped back into video domain using our pre-trained 3D-MBQ-VAE decoder $\mathcal{D}$.
@@ 5 Experimental results
We now present experimental results. The details of experimental setup are provided in Appendices C and D. Additional qualitative results are provided in Appendices A, G and F.
@@ 5.1 Latent Reconstruction Results of 3D-MBQ-VAE
For evaluating our 3D-MBQ-VAE, we selected COCO-2017 and WebVID validation datasets. Following Zhao et al. (2024), we crop each frame to 256$\times$256 resolution and sample 48 frames per video sequentially. As shown in Table 1, our proposed 3D-MBQ-VAE consistently outperforms SOTA 3D VAEs across all metrics. This can be attributed to a novel combination of training strategies and the enhanced loss function. Randomly masking parts of various frames He et al. (2021) enables our model to efficiently learn spatial information. Further, inclusion of $\mathcal{L}_{\text{MFI}}$ in the loss function emphasizes learning efficient temporal features. On removing $L_{\text{mfi}}$, LPIPS metric decreases substantially. This highlights its role in preserving perceptual quality. Our 3D-MBQ-VAE model is also capable of supporting joint image and video training, similar to W.A.L.T Gupta et al. (2025) and CV-VAE. As shown in Table 1, joint training provides superior results than video training alone.
Table 1: Quantitative comparison of VAEs for video reconstruction task on COCO-Val and WebVid-Val datasets. Frame compression rate (FCR) is the ratio between the size of video frame before and after compression. Compatibility (comp.) represents if the model can be used as a VAE for existing generative models. Our method demonstrates superior performance on all metrics. All models are trained on videos, except the “3D-MBQ-VAE (Video+Images)” variant that is jointly trained on video and images. Hence, its results are not compared with other methods.
@@
3D-MBQ-VAE (Without $L_{\text{mfi}}$) 140M + 177M 4$\times$ 30.8 / 0.840 / 0.112 31.4 / 0.849 / 0.134
3D-MBQ-VAE 140M + 177M 4$\times$ 31.2 / 0.848 / 0.092 32.1 / 0.858 / 0.108
3D-MBQ-VAE (Video+Images) 140M + 177M 4$\times$ 33.0 / 0.866 / 0.087 34.2 / 0.877 / 0.092
@@
Figure 8: The videos corresponding to these frames show the improvement in motion quality brought by FFT layers. For videos, click on link.
@@ 5.2 Results of Text Conditioned Video Generation
Table 3: Results of the text-conditioned video generation (Text2Video) models. For all the techniques, evaluation was done on MSR-VTT dataset. Recaptioning the dataset improves all the metrics. Inference Time was calculated on a single A100.
Method Params Train: WebVID 10M WebVid 10M - recaptioned Inf Tims(s)
@@ 6 Conclusion
Acknowledgement: Support for this work was provided by Science and Engineering Research Board (SERB) of India, under the project CRG/2022/003821. Jishu and Chirag contributed to this project while working as interns at IIT Roorkee.
@@
Encoder The encoder takes an input video with dimensions $B\times C\times N\times H\times W$, where $B$ is the batch size, $C$ is the number of channels, and $N\times H\times W$ represents the spatiotemporal dimensions (i.e., frames, height, and width). The encoding begins with a Conv3D block that captures spatiotemporal features using a 3D convolution with a kernel size of $1\times 3\times 3$, followed by a stride of $1\times 2\times 2$ to reduce spatial resolution while expanding the channel depth (e.g., from 3 to 64 channels). The encoder then applies multiple 3D-MB Conv blocks (Mobile Inverted Conv Blocks) to capture complex features using depthwise separable convolutions, maintaining computational efficiency. These blocks are interspersed with Inception-Fused-MB Conv blocks, which improve feature extraction by combining multiple filter sizes and dimensions, enabling multi-scale processing.
@@ B.1.2 Key Features
B.1.3 Mobile-Inverted 3D Conv Block
@@
B.1.4 Inception-Fused MB Conv Block
@@
The 3D Inception-Fused-MB Conv Block is a sophisticated convolutional module designed to efficiently capture both spatial and temporal features in video data, commonly used in tasks like Video VQ-VAE (Vector Quantized Variational Autoencoder) shown in Figure S.8. This block employs an Inception-style structure with three parallel 3D convolution paths, each with different kernel configurations (e.g., 3 x 3 x 1, 3 x 1 x 3, and 1 x 3 x 3) to capture multi-scale spatiotemporal patterns in various dimensions. Each branch is followed by a depthwise separable 3D-MobileBlock (MB-Conv) that reduces computational complexity while maintaining efficiency. The outputs of these paths are concatenated, allowing the network to fuse multi-scale features, which are then processed by a final 3D convolutional block. This design enables efficient handling of complex video representations in VQ-VAEs, optimizing the model’s ability to compress and reconstruct video sequences while preserving key spatial and temporal information. We perform temporal downsampling within the 3D-MBQ-VAE encoder at specific stages. Temporal downsampling is applied by a factor of 2 at the 6th block (Downsampling Block) and again by a factor of 2 at the 9th block, resulting in a total downsampling factor of 4 along the temporal dimension.
B.2 Additional Abblation Studies for 3D-MBQ VAE
B.2.1 Impact of varying codebook sizes on video reconstruction
We conducted a comprehensive ablation study analyzing the effects of different codebook sizes, embedding dimensions, and quantization techniques on video reconstruction quality. In our default configuration, we employ the Lookup-Free Quantization (LFQ) method from MAGVIT-v2, utilizing an embedding size of $d=8$ and a codebook vocabulary size of $C=12,800$.
Table 7: Impact of codebook size ($C$) on video reconstruction (d= Embedding Size)
COCO-Val WebVid-Val
Method Of Quantization $C$ $d$ Time (ms) PSNR/ SSIM/ LPIPS PSNR/ SSIM/ LPIPS
VQ 256 512 35 29.96/ 82.21/ 0.133 30.01/ 81.89/ 0.1441
Gumbal Quantization 256 512 23 29.83/ 81.98/ 0.140 29.78/ 80.56/ 0.1511
VQ 1024 256 47 30.26/ 82.54/ 0.130 30.72/ 81.09/ 0.1453
VQ 4096 256 56 30.72/ 83.11/ 0.127 31.03/ 82.22/ 0.1331
VQ 8000 128 69 30.45/ 82.19/ 0.144 30.98/ 83.36/ 0.1299
LFQ 8000 16 20 31.08/ 84.34/ 0.112 31.98/ 85.04/ 0.1103
LFQ(Default) 12800 8 22 31.22/ 84.78/ 0.092 32.09/ 85.78/ 0.1081
The primary reasons for adopting LFQ over traditional Vector Quantization (VQ) methods are:
1. Larger Vocabulary with Smaller Embedding Size: LFQ enables the use of a considerably larger codebook vocabulary while maintaining a compact embedding size. This approach enhances the model’s expressive power without significantly increasing computational costs. By leveraging a larger vocabulary with smaller embeddings, the model benefits from higher codebook utilization, a strategy that has been shown to improve performance Sun et al. (2024c) for the task of image reconstruction.
2. Improved Efficiency: LFQ is faster and more optimized than standard VQ methods, which is beneficial for processing high-resolution video data.
B.2.2 Results with random masking and full-frame masking
For random spatial masking, we employ a “cosine scheduling” strategy, progressively varying the masking ratio from 20% to 60% during training. This scheduling helps the model gradually adapt to different levels of partial observations, thereby enhancing robustness in spatial feature extraction.
For frame masking, we similarly utilize cosine scheduling, adjusting the masking ratio from 10% to 50%. This approach ensures a balanced training dynamic where the model is exposed to varying levels of temporal information occlusion, ultimately aiding in better temporal coherence and reconstruction quality. These masking strategies are designed to incrementally challenge the model, helping it learn to reconstruct meaningful content under different masking conditions and thus improving overall generalization.
Table 8 shows the results with random masking and full-frame masking. These results underscore the significance of both random masking and fully masked frames in enhancing the model’s ability to predict frame indices effectively.
Table 8: Ablation results with random masking and full-frame masking
COCO-Val WebVid-Val
Methods PSNR/ SSIM/ LPIPS PSNR/ SSIM/ LPIPS
W/o Random Masking 29.17/ 81.92/ 0.111 28.78/ 82.29/ 0.1334
W/o Full Frame Masking 30.09/ 82.22/ 0.102 30.11/ 83.01/ 0.1404
W/o Random Masing and Full frame Masking 28.82/ 80.77/ 0.124 26.77/ 79.92/ 0.1552
With Random Masing and Full frame Masking 31.22/ 84.78/ 0.092 32.09/ 85.78/ 0.1081
The random masking strategy compels the model to infer missing spatial information from partial observations within each frame, effectively enhancing its ability to capture spatial dependencies and structures. By being trained on randomly masked data, the model learns to reconstruct or predict spatial details based on contextual cues, leading to stronger spatial representations. Conversely, supervision for fully masked frame index prediction requires the model to determine the correct temporal order of completely obscured frames, which enhances temporal consistency. This task forces the model to understand and model temporal dynamics across frames, as it must rely on learned temporal patterns to accurately predict the positions of fully masked frames within the sequence.
B.2.3 Video compression results on MCL-JCV dataset
We conducted a zero-shot inference on 30 videos from the MCL-JCV dataset (Wang et al. (2016)), resized to a resolution of 640 x 360, which aligns with the experimental setup used by MAGVIT Yu et al. (2023a) and MAGVIT-v2Yu et al. (2023c) . We evaluate compression quality using standard distortion metrics (LPIPS, PSNR, and SSIM) at a bit rate of 0.0384 bpp (bits per pixel). The results are shown in Table 9. At equivalent bit rates, traditional codecs such as HEVC Sullivan et al. (2012) and VCC Bross et al. (2021) may sometimes achieve finer local detail rendering compared to 3D-MBQ-VAE. However, they often introduce block artifacts that, while detrimental to perceptual quality, are not adequately captured by PSNR and SSIM. This is reflected in the LPIPS metric.
Table 9: Video compression results on MCL-JCV dataset
Methods PSNR SSIM LPIPS
HEVC 30.1 0.943 0.199
VCC 32.65 0.966 0.153
MAGVIT 23.7 0.846 0.144
MAGVIT-v2 26.18 0.894 0.104
3D-MBQ-VAE (Ours) 29.09 0.922 0.089